# Quantitative assessment of biological dynamics with aggregate data

### v1

@@ Abstract
We develop and apply a learning framework for parameter estimation in initial value problems that are assessed only indirectly via aggregate data such as sample means and/or variances. Our comprehensive framework follows Bayesian principles and consists of specialized Markov chain Monte Carlo computational schemes that rely on modified Hamiltonian Monte Carlo to align with summary statistic constraints and a novel elliptical slice sampler adapted to the parameters of biological models. We benchmark our methods with synthetic data on microbial growth in batch culture and test them on real growth curve data from laboratory replication experiments on Prochlorococcus microbes. The results indicate that our learning framework can utilize experimental or historical data and lead to robust parameter estimation and data assimilation in ODE models of biological dynamics that outperform least-squares fitting.
@@ 1 Introduction
The practice of acquiring large sets of individual data points and combining them to obtain diverse summary statistics, which we refer to as data aggregation, is a prevalent technique used in multiple domains, including economics, policy making, social sciences, healthcare, and biological and ecological research [26, 63, 40]. In modern science and engineering, the interpretation of aggregate data is often preferable to the interpretation of raw data, and this practice is represented in all high-level data analysis contexts.
@@
Specifically, in the life and biological sciences aggregate data is often generated in replication experiments where the experiment’s actual raw measurements are used to generate summary statistics, such as averages and standard errors, that are maintained, curated, and made openly available while the original raw measurements are either discarded or access to them is kept restricted. This practice has the obvious advantages of conserving storage space when processing large data sets such as time-lapse and image data [31], obscuring the actual source of the data for security and privacy or competitive purposes [66, 20], reducing noise [68, 16], and aiding comparison between different experiments or protocols that are probing the same system but employing different modalities [67]. However, such practice leads to data collection problems, for example having only summary statistics and having difficulty acquiring raw data for a complete understanding of Covid-19 data affected maternal and neonatal outcomes [59]. An additional challenge, perhaps more significant, of working with pre-processed data is the loss of information that occurs when generating it [49] and also the loss of reference to the underlying specific biological processes [56].
@@
Data aggregation also poses serious challenges with rigorous data analysis or assimilation techniques [30]. Specifically, within data assimilation, the goal is to develop mathematical frameworks that process information available in empirical form and obtain quantitative predictions [65]. However, predictions following the assimilation of aggregated data are limited by lost information or distortions caused by aggregation [22, 15]. In order to counteract such artifacts, elaborate frameworks that combine domain knowledge and physical constraints in the form of specialized models are often required. Such approaches naturally lend themselves to Bayesian assimilation methodologies [53].
@@
Bayesian data assimilation methods have seen an increase in popularity, particularly in parameter estimation applications for dynamical problems modeled with initial value problems (IVP) [24, 23, 36]. A major advantage is that they offer uncertainty quantification, as the standard practice is to learn an entire distribution of plausible values for every variable of interest rather than point estimators. Another advantage is that they can restrict the parameters under estimation to only meaningful values; for instance, a rate parameter or an initial population can be assigned a prior distribution with only positive support, or prior distributions can be defined on only the intervals that remain meaningful in the scales of the underlying problem. Furthermore, they distinguish identifiable and nonidentifiable cases, the distinction of which is a common modeling challenge in mathematical biology [38, 64]. Finally, they offer modeling flexibility and realism that comes from explicitly accounting for different sources of noise and complexity characteristic of the systems being studied [51].
@@
Nevertheless, Bayesian data assimilation is not without its drawbacks, especially for IVPs. To start, one needs to construct and characterize, often via intensive computational sampling procedures, the relevant posterior distribution. In particular, in the parameter estimation problem of ODE models, this involves numerically integrating the ODE for each new configuration of parameters for each sample [46, 55, 1]. Given that required sample sizes are typically large, sampling strategies become computationally costly and require efficient computational algorithms that are the topic of ongoing research. In addition, biological dynamics are typically contaminated with multiplicative noise [12, 13, 23] which poses additional challenges to many existing algorithms which assume additive noise. Furthermore, the predictions provided in Bayesian data assimilation depends critically on the model used, its fidelity to the modeled system, and the quality of the data supplied.
@@
A particular challenge when training a dynamical model with aggregated data is that, to properly model the data points that give rise to the supplied summary statistics, we have to mathematically reproduce the collapsing of unavailable raw measurements down to the available values. In the statistical representation of the resulting model, this translates into fitting the model parameters with singular distributions, that is, probability distributions supported only on subspaces of lower dimension than the model’s full parameter space. Parameter estimation is a routine topic in the Bayesian literature focused on simple models with nonsingular distributions [58, 51]; however, parameter estimation with singular distributions under complex statistical models necessary to tackle real-world biological scenarios remains an open challenge.
@@
In this study, we develop a novel comprehensive statistical learning framework that addresses these challenges. Our framework allows for parameter estimation in IVPs of ODE models and also allows modeling of the latent raw measurements that gave rise to the summary statistics forming our dataset. To perform parameter estimation on our statistical model, we apply methods that can sample from distributions restricted to particular subsets of the parameter space as determined by data aggregation. To this end, we develop a novel extension of the Hamiltonian Monte Carlo (HMC) sampling algorithm that allows for navigating highly dimensional parameter spaces while accounting for parameter constraints.
@@
The rest of this study is structured as follows. In Methods, we begin by describing the data set under consideration, the IVP we aim to address, and the Bayesian framework used for parameter estimation. First, we present the Bayesian framework in general form, followed by its application to the analysis of growth curve data. In Results, we demonstrate how our framework performs and compare it with in silico and in vitro data of microbial growth curves from batch culture experiments commonly conducted in microbiology. Lastly, in Discussion, we provide an overview of our methods and elaborate on its perspectives for future applications.
@@ 2 Methods
In this section, we first present a general description of our framework that emphasizes modeling and computational aspects. Subsequently, we present a specialized application to microbial growth curves obtained in replication experiments.
@@ 2.1.1 Learning scheme
Here, $y_{n}^{k}$ denote individual raw measurements, made at time $t_{n}$, and $k=1,\dots,K$ indexes the batch of measurements. The functions $G^{m}(\cdot)$ model the corresponding batch statistics. For instance, the common sample mean and variance correspond to the functions
@@
$G^{2}\left(y^{1:K}\right)$ $=\frac{1}{K-1}\sum_{k=1}^{K}\left(y^{k}-\frac{1}{K}\sum_{k^{\prime}=1}^{K}y^{k^{\prime}}\right)^{2}.$ (3)
In this study, we focus on the cases where only $z_{1:N}^{1:M}$ are available, while $y_{1:N}^{1:K}$ are not. For this reason, we explicitly model the missing measurements using a likelihood that represents biological or measurement noise. Our likelihood
$y_{n}^{k}|g,h\sim\mathbb{A}\left(F\left(x^{g}(t_{n})\right),h\right)$ (4)
@@
The function dynamics $H^{g}(\cdot,\cdot)$ is also problem specific and depends on unknown parameters $g$. Together, with the appropriate initial conditions that may also depend on unknown parameters $g$, our IVP leads to a solution $x^{g}(t)$ which, at any time $t$, describes the dynamical state of the system of interest.
Our framework contains unknown measurements $y_{1:N}^{1:K}$ whose statistical properties are fully described by the likelihood in eq. 4 and also unknown parameters $g$ and $h$. The latter affects the noise of the system, while the former affects the dynamics. Following the Bayesian paradigm [51], we assign priors to them
@@
which, besides completing our statistical framework, allow the specification of meaningful bounds for their values.
Figure 1: Graphical representation of the statistical learning framework in this study. Following the common convention, random quantities are shown with circles, deterministic quantities are shown with diamonds, and quantities with known values are shaded. In addition, arrows indicate dependencies among the various quantities of interest, plates indicate repetition over the marked index, and hyperparameters are left free.
@@
Here, the factors containing Dirac deltas $\delta_{G^{m}\left(\cdot\right)}\left(\cdot\right)$ arise because of eq. 1, which dictate precise agreement between our model measurments and the corresponding batch statistics, while $A(\cdot|\cdot,\cdot),B(\cdot),C(\cdot)$ are the probability densities associated with the distributions $\mathbb{A}(\cdot,\cdot),\mathbb{B},\mathbb{C}$.
@@ 2.1.2 Markov chain Monte Carlo
Due to the IVP, which most often remains analytically intractable, we generally cannot obtain a closed-form expression to the posterior of eq. 8. For this reason, to characterize eq. 8, we develop a specialized Markov chain Monte Carlo (MCMC) sampling scheme that generates pseudorandom samples of the model unknowns $h,g,y_{1:N}^{1:K}$ [37, 54, 42].
Because of the natural grouping of our unknowns, we use a Gibbs sampler that iterates the following two steps:
@@
To initialize our sampler, we directly sample $h$ and $g$ from the priors $\mathbb{B}$ and $\mathbb{C}$, respectively. Next, we sample $y_{n}^{k}\mid g,h,$ from $\mathbb{A}$, and then apply root-finding on the generated $y_{1:N}^{1:M}$ to ensure that the constraints are satisfied. In all cases, we combine our sampler with appropriate numerical integrators for the solution of the IVP [52, 60, 3, 57] adapted to the specifics of the problem at hand.
In the special case where $\mathbb{A}$ and $\mathbb{B}$ are conditionally conjugate, our Gibbs updates can be implemented via ancestral sampling [7, 37, 18, 51] based on the factorizations
@@
both of which allow for direct sampling of $h$ via $\mathcal{P}\left(h|g,y_{1:N}^{1:K}\right)$. For both the parameters and the measurements, our updates are derived from specialized samplers adapted to $\mathcal{P}\left(g|y_{1:N}^{1:K}\right)$ and $\mathcal{P}\left(y_{1:N}^{1:K}|g,z_{1:N}^{1:M}\right)$ as described below.
@@
The first Gibbs update requires the sampling of $g$ from $\mathcal{P}\left(g|y_{1:N}^{1:K}\right)$, which we achieve using a novel multiplicative elliptical slice sampler (mESS), which naturally aligns with distributions restricted to positive support as often found in dynamical systems of biological processes. See section C.1 for a detailed explanation of the sampler. Our approach retains the core benefits of elliptical slice sampling, such as parameter updates without tuning and efficient exploration of complex posterior landscapes, while offering a targeted enhancement for distributions with strictly positive values and non-Gaussian architecture [51, 47].
@@
In our second Gibbs update, we generate samples of $y_{1:N}^{1:K}$ given known parameters $g$ and $h$ that satisfy the constraints $z_{n}^{m}=G^{m}(y_{n}^{1:K})$ of eq. 1. For this, we apply a novel constrained Hamiltonian Monte Carlo (cHMC) method specialized for handling the constraints. Our method is fully detailed in section C.3. As we explain in section C.2, standard HMC is suitable for sampling smooth high-dimensional distributions with full support [37, 8, 5]; however, satisfying the constraints requires modifications. Our novel cHMC sampler takes advantage of the RATTLE integrator [2, 9, 21] in the HMC integration loop to ensure that the Monte Carlo samples of $y_{1:N}^{1:K}$ satisfy the constraints while remaining statistically correct, i.e. maintaining that the MCMC chain converges to the target distribution. Our method maintains HMC’s efficient sampling of high-dimensional distributions [37, 8, 5] arising due to multiple time points and large batch-sizes, while permitting navigation on the support of singular distributions arising due to the constraints.
@@ 2.2 Application to Prochlorococcus growth curve data
In this section, we specialize our statistical learning framework in a case study of interest in microbiology and marine biology. In our study, the aggregated data stem from batch growth experiments of the marine cyanobacteria Prochlorococcus (Pro). Pro is the most abundant photosynthetic phytoplankton in the ocean [50] and is studied in situ for genetic and physical connections to their bio-geographical significance [17, 4]. These photosynthetic organisms play a vital role in the regulation of ocean food chains and climate [6, 50]. Laboratory growth experiments take the form of replicate time series data, where multiple sets of triplicate test tubes of Pro are monitored for cell density as they grow in batch culture. Data aggregation is applied to the runs to produce sample averages and variances.
Our growth curves depend on two pivotal quantities: the maximum growth rate, typically reported in units of $1/\text{\rm days}$, and the nutrient affinity, typically reported in units $m^{3}/(\text{\rm cell}\cdot\text{\rm days})$, of the Pro cells which we denote with $m$ and $a$, respectively. These parameters, as well as the initial nutrient and Pro cell densities of batch culture experiments, which we denote with $Q$ and $P$, determine the overall dynamics of the cell and nutrient density.
@@
$\frac{dq}{dt}$ $=-\frac{q}{q+\frac{m}{a}}mp$ $q(t_{0})$ $=Q$ (9)
$\frac{dp}{dt}$ $=+\frac{q}{q+\frac{m}{a}}mp$ $p(t_{0})$ $=P$ (10)
for some fixed initial time $t_{0}$. Our dynamical state $x=(q,p)$ consists of the density of nutrients $q$ and the density of cells $p$, both reported in $\text{\rm cells}/m^{3}$. Our eqs. 9 and 10 depend on the unknown parameters $g=(P,Q,m,a)$. Given $g$, we denote the solution of our IVP with $x^{g}(t)=(q^{g}(t),p^{g}(t))$.
@@
Theoretical and empirical studies on microbial growth [12, 13, 23] indicate the presence of multiplicative noise in the measurements. Accordingly for the likelihood $\mathbb{A}(\cdot,\cdot)$ we choose the LogNormal distribution
$y_{n}^{k}|g,h\sim\text{LogNormal}\left(p^{g}(t_{n}),h\right)$ (12)
with an unknown scale parameter $h$. For the definition, see appendix B.
Following section 2.1.1, we denote by $y_{n}^{k}$ the $k^{\text{th}}$ measurement made at time $t_{n}$ during the experiment and assume the statistics in eqs. 2 and 3 to form the reported mean $z^{1}_{n}$ and variance $z^{2}_{n}$ of each time point.
Finally, for eq. 6, we apply a prior on the parameters $g=(P,Q,m,a)$ of our IVP that allows only for strictly positive values as defined by
$Q$ $\sim\text{Gamma}\left(\phi_{Q},\frac{\psi_{Q}}{\phi_{Q}}\right),$ (13)
$P$ $\sim\text{Gamma}\left(\phi_{P},\frac{\psi_{P}}{\phi_{P}}\right),$ (14)
$m$ $\sim\text{Gamma}\left(\phi_{m},\frac{\psi_{m}}{\phi_{m}}\right),$ (15)
$a$ $\sim\text{Gamma}\left(\phi_{a},\frac{\psi_{a}}{\phi_{a}}\right).$ (16)
@@
$h\sim\text{Gamma}\left(\phi_{h},\frac{\psi_{h}}{\phi_{h}}\right),$ (17)
which is conditionally conjugate to the likelihood in eq. 12. Here we use the parameterization of the Gamma distribution shown in appendix B.
Our framework leads to the derivation of a posterior density
@@
here we marginalize the noise parameter $h$ since its value is of little biological interest. By this marginalization, we also avoid the ancestral sampling step in our MCMC implementation of section 2.1.2.

### v2

@@ Abstract
We develop and apply a learning framework for parameter estimation in initial value problems that are assessed only indirectly via aggregate data such as sample means and/or standard deviations. Our comprehensive framework follows Bayesian principles and consists of specialized Markov chain Monte Carlo computational schemes that rely on modified Hamiltonian Monte Carlo to align with constraints induced by summary statistics and a novel elliptical slice sampler adapted to the parameters of biological models. We benchmark our methods with synthetic data on microbial growth in batch culture and test them with real growth curve data from laboratory replication experiments on Prochlorococcus microbes. The results indicate that our learning framework can utilize experimental or historical data and lead to robust parameter estimation and data assimilation in ODE models that outperform least-squares fitting.
@@ 1 Introduction
The practice of acquiring large sets of individual data points and combining them to obtain diverse summary statistics, which we refer to as data aggregation, is a commonly used technique in multiple domains, including economics, policy making, social sciences, health care and biological and ecological research [30, 85, 48]. In modern science and engineering, the interpretation of aggregate data is often preferable to the interpretation of raw data, and this practice is represented in all high-level data analysis contexts.
@@
Specifically, in the life and biological sciences aggregate data is often generated in replication experiments where the experiment’s actual raw measurements are used to generate summary statistics, such as averages and standard errors, that are maintained, curated, and made openly available while the original raw measurements are either discarded or access to them is kept restricted. This practice has the obvious advantages of conserving storage space when processing large data sets such as time-lapse and image data [39], obscuring the actual source of the data for security and privacy or competitive purposes [88, 23], reducing noise [90, 17, 76, 75, 73], and aiding comparison between different experiments or protocols that are probing the same system but employing different modalities [89, 77, 73, 74]. However, such practices can cause problems; for example, many Covid-19 data related to maternal and neonatal outcomes are restricted to only summary statistics, leaving clinicians and patients to operate on partial information [80].
An additional challenge, perhaps more significant, of working with aggregate data is the loss of information that occurs when it is generated [59] and also the loss of reference to the underlying specific biological processes [68, 63]. For example, similar to averaging, the integration of signals that occurs during the acquisition of single-molecule fluorescence data obscures fast dynamics without allowing for the precise estimation of detailed kinetics [33, 35, 32, 34, 50]. In addition, biological specimens often exhibit heterogeneity, and aggregating data without accounting for heterogeneity may produce skewed results [72, 71]. Accurate quantitative analysis requires careful experimental design and data processing to avoid masking subtle differences between individual specimens. Under these conditions, reproducibility is particularly challenging, as variability in biological systems can make replicated findings difficult due to small differences that remain uncharacterized prior to data acquisition.
@@
Data aggregation also poses serious challenges with rigorous data analysis or assimilation techniques [38]. Specifically, within data assimilation, the goal is to develop mathematical frameworks that process the information available in empirical form and obtain quantitative predictions [87]. However, predictions following the assimilation of aggregated data are limited by lost information or distortions caused by aggregation [25, 16]. In order to counteract such artifacts, elaborate frameworks that combine domain knowledge and physical constraints in the form of specialized models are often required. Such approaches naturally lend themselves to Bayesian assimilation methodologies [64].
@@
Bayesian data assimilation methods have seen an increase in popularity, particularly in parameter estimation applications for dynamical problems modeled with initial value problems (IVP) [27, 26, 44]. A major advantage is that they offer uncertainty quantification, as the standard practice is to learn an entire distribution of plausible values for every variable of interest rather than point estimators. Another advantage is that they can restrict the parameters under estimation to only meaningful values; for instance, a rate parameter or an initial population can be assigned a prior distribution with only positive support, or prior distributions can be defined on only the intervals that remain meaningful in the spatiotemporal scales of the underlying problem. Furthermore, they apply equally to both identifiable and nonidentifiable cases, the distinction of which is a common modeling challenge in mathematical biology [46, 86]. Finally, they offer modeling flexibility and realism that comes from explicitly accounting for different sources of noise and the complexity characteristic of the systems being studied [61].
@@
Nevertheless, Bayesian data assimilation is not without drawbacks, especially for IVPs [67, 70]. To start, one needs to construct and characterize, often via intensive computational sampling procedures, the relevant posterior distribution. In particular, in the parameter estimation problem of models with ordinary differential equations (ODE), this involves numerically integrating the ODE for each new configuration of parameters for each posterior sample [56, 66, 1]. Given that the required posterior sample sizes are typically large, sampling strategies become computationally costly and require efficient computational algorithms that are a topic of ongoing research. In addition, biological dynamics are typically contaminated with multiplicative noise [13, 14, 26], which poses additional challenges to common algorithms that assume additive noise [83, 20], such as sudden divergence of Kalman filters [61, 8] or degeneracy of particle filters [18]. Furthermore, the predictions provided in Bayesian data assimilation depends critically on the dynamical model used, its fidelity to the modeled system, and the quality of the data supplied.
@@
A particular challenge when training a dynamical model with aggregated data is that, to properly model the data points that give rise to the supplied summary statistics, we have to mathematically reproduce the collapsing of unavailable raw measurements down to the available values. In the statistical representation of the resulting model, this translates into fitting the model parameters with singular distributions, that is, probability distributions supported only on subspaces of lower dimension than the model’s full parameter space. Parameter estimation is a routine topic in the Bayesian literature focused on simple models with nonsingular distributions [79, 61]; however, parameter estimation with singular distributions under complex statistical models necessary to tackle real-world scenarios remains an open challenge.
@@
In this study, we develop a novel comprehensive statistical learning framework that addresses these challenges. Our framework allows for parameter estimation in IVPs of ODE models and also allows modeling of the latent raw measurements that gave rise to the summary statistics forming our dataset. To perform parameter estimation on our statistical model, we apply computational methods that can sample from distributions restricted to particular subsets of the parameter space determined by data aggregation. To this end, we develop a novel extension of the Hamiltonian Monte Carlo (HMC) sampling algorithm that allows for navigating highly dimensional parameter spaces while accounting for parameter constraints. In addition, we develop a novel slice sampling scheme that allows for the IVP parameter’s positive support to be exploited efficiently by means of the elliptical slice sampling (ESS) algorithm.
@@
The rest of this study is structured as follows. In Methods, we begin by describing the data format under consideration, the IVP we aim to address, and the Bayesian framework that we use for its parameter estimation. First, we present our framework in general form, followed by its application to the analysis of growth curve data. In Results, we demonstrate how our framework performs and compares it with standard estimation approaches, such as least squares fitting, with in silico and in vivo data of microbial growth curves from batch culture laboratory experiments commonly conducted in microbiology research. Lastly, in Discussion, we provide an overview of our methods and elaborate on its perspectives for future applications.
@@ 2 Methods
In this section, we first present a general description of our framework that emphasizes modeling and computational aspects. Subsequently, we present a specialized application to microbial growth curve data acquired in replication experiments.
@@ 2.1.1 Learning scheme
Here, $y_{n}^{k}$ denotes individual raw measurements, made at time $t_{n}$, and $k=1,\dots,K$ indexes the batch of measurements. The functions $G^{m}(\cdot)$ model the corresponding batch statistics. For instance, the common sample mean and standard deviation correspond to the functions
@@
$G^{2}\left(y^{1:K}\right)$ $=\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(y^{k}-\frac{1}{K}\sum_{k^{\prime}=1}^{K}y^{k^{\prime}}\right)^{2}}.$ (3)
In this study, we focus on the cases where only $z_{1:N}^{1:M}$ are available, while $y_{1:N}^{1:K}$ are not. For this reason, we explicitly model the missing measurements using a likelihood that represents biological or measurement noise. Our likelihood, which we cast in the form
$y_{n}^{k}|g,h\sim\mathbb{A}\left(F\left(x^{g}(t_{n})\right),h\right),$ (4)
@@
The dynamics function $H^{g}(\cdot,\cdot)$ is also problem specific and depends on unknown parameters gathered in $g$. Together, with the appropriate initial conditions that may also depend on unknown parameters gathered in $g$, our IVP leads to a solution $x^{g}(t)$ which, at any time $t$, describes the dynamical state of the system of interest.
Our framework contains unknown measurements $y_{1:N}^{1:K}$ whose statistical properties are fully described by the likelihood in eq. 4 and also unknown parameters $g$ and $h$. The latter is associated with the noise of the system, while the former is associated with the dynamics. Following the Bayesian paradigm [61], we assign priors to them
@@
which, in addition to completing our statistical framework, allow the specification of meaningful bounds for their values through their support.
Figure 1: Graphical representation of the statistical learning framework in this study. Following the common convention [61, 7], random quantities are shown with circles, deterministic quantities are shown with diamonds, and quantities with known values are shaded. In addition, arrows indicate dependencies among the various quantities of interest, plates indicate repetition over the marked index, and hyperparameters are left free.
@@
Here, the factors that contain Dirac deltas $\delta_{G^{m}\left(\cdot\right)}\left(\cdot\right)$ arise due to eq. 1, which dictates the precise agreement between our model measurements and the corresponding batch statistics, while $A(\cdot|\cdot,\cdot),B(\cdot),C(\cdot)$ are the probability functions associated with the distributions $\mathbb{A}(\cdot,\cdot),\mathbb{B},\mathbb{C}$ in eqs. 4, 6 and 7, respectively.
@@ 2.1.2 Markov chain Monte Carlo
Due to the IVP, which most often remains analytically intractable, we generally cannot obtain a closed-form expression to the posterior of eq. 8. For this reason, to characterize eq. 8, we develop a specialized Markov chain Monte Carlo (MCMC) sampling scheme that generates pseudorandom samples [45, 65, 52] of the unknowns of the model $h,g,y_{1:N}^{1:K}$.
Due to the natural grouping of our unknowns, we use a Gibbs sampler that iterates the following two steps:
@@
To initialize our sampler, we directly generate $h$ and $g$ from the priors $\mathbb{B}$ and $\mathbb{C}$, respectively. Next, we generate $y_{n}^{k}\mid g,h,$ from $\mathbb{A}$, and then apply root-finding on the generated $y_{1:N}^{1:M}$ to ensure that the constraints of eq. 1 are satisfied. In all cases, we combine our sampler with appropriate numerical integrators, such as adaptive Runge-Kutta or other specialized schemes, to solve the IVP according to the specifics of the problem at hand [62, 81, 3, 78].
In the special case where $\mathbb{A}$ and $\mathbb{B}$ are conditionally conjugate [21], our Gibbs updates can be implemented via ancestral sampling [7, 45, 21, 61] based on the respective factorizations
@@
both of which allow direct generation of $h$ via $\mathcal{P}\left(h|g,y_{1:N}^{1:K}\right)$. For both the parameters and the measurements, our updates are derived from specialized samplers adapted to $\mathcal{P}\left(g|y_{1:N}^{1:K}\right)$ and $\mathcal{P}\left(y_{1:N}^{1:K}|g,z_{1:N}^{1:M}\right)$ as described below.
@@
The first Gibbs update requires the generation of $g$ given measurements $y^{1:K}_{1:N}$ from $\mathcal{P}\left(g|y_{1:N}^{1:K}\right)$, which we achieve using a novel multiplicative elliptical slice sampler (mESS), which naturally aligns with distributions restricted to positive support, as often found in dynamical systems of biological processes. See section C.1 for a detailed explanation of the sampler. Our approach retains the core benefits of elliptical slice sampling, such as parameter updates without tuning and efficient exploration of complex posterior landscapes, while offering a targeted enhancement for distributions with strictly positive values and non-Gaussian characteristics [61, 57].
@@
In our second Gibbs update, we generate samples of $y_{1:N}^{1:K}$ given known parameters $g$ from $\mathcal{P}\left(y_{1:N}^{1:K}|g,z_{1:N}^{1:M}\right)$ that satisfy the constraints $z_{n}^{m}=G^{m}(y_{n}^{1:K})$ of eq. 1. For this, we apply a novel constrained Hamiltonian Monte Carlo (cHMC) sampler specialized for handling the constraints. Our method is fully detailed in section C.3. As we explain in section C.2, standard HMC is suitable for sampling smooth high-dimensional distributions with full support [45, 9, 5]; however, satisfying the constraints requires modifications as described in section C.3. Our novel cHMC sampler takes advantage of the RATTLE numerical integrator [2, 10, 24] in the HMC integration loop to ensure that the generated $y_{1:N}^{1:K}$ satisfy the constraints while remaining statistically correct, i.e. ensuring that our MCMC chain converges to the target distribution of eq. 8. Our method maintains HMC’s efficient sampling of high-dimensional distributions [45, 9, 5] arising due to multiple time points (i.e. $N\gg 1$) and large batch-sizes (i.e. $K\gg 1$) per time point, while permitting navigation on the support of singular distributions arising due to the constraints.
@@ 2.2 Application to Prochlorococcus growth curve data
In this section, we specialize our statistical learning framework in a case study of interest in microbiology and marine biology. In our study, the aggregated data stem from batch growth experiments of the marine cyanobacteria Prochlorococcus (Pro). Pro is the most abundant photosynthetic phytoplankton in the ocean [60] and is studied in situ for genetic and physical connections to their bio-geographical significance [19, 4]. These photosynthetic organisms play a vital role in the regulation of ocean food chains and climate [6, 60]. Laboratory growth experiments take the form of replicate time series data, where multiple sets of triplicate test tubes of Pro are monitored for cell density as they grow in batch culture. Data aggregation is applied to the runs to produce sample averages and standard deviations.
Our growth curves depend on two pivotal quantities: the maximum growth rate, typically reported in units of $1/\text{\rm days}$, and the nutrient affinity, typically reported in units $\text{\rm ml}/(\text{\rm cells}\cdot\text{\rm days})$, of the Pro cells which we denote with $m$ and $a$, respectively. These parameters, as well as the initial nutrient and Pro cell densities of batch culture experiments, which we denote with $Q$ and $P$, determine the overall dynamics of the cell and nutrient density.
@@
$\frac{dq}{dt}$ $=-\frac{q}{q+m/a}mp$ $q(t_{0})$ $=Q$ (9)
$\frac{dp}{dt}$ $=+\frac{q}{q+m/a}mp$ $p(t_{0})$ $=P$ (10)
for a fixed initial time $t_{0}$ that coincides with the onset of the experiment. Our dynamical state $x=(q,p)$ consists of the density of nutrients $q$ and the density of cells $p$, both reported in cells/ml. Our eqs. 9 and 10 depend on the unknown parameters $g=(Q,P,m,a)$. Given $g$, we denote the solution of our IVP with $x^{g}(t)=(q^{g}(t),p^{g}(t))$.
@@
Theoretical and empirical studies on microbial growth [13, 14, 26] indicate the presence of multiplicative noise in the measurements. Accordingly, for the likelihood $\mathbb{A}(\cdot,\cdot)$ we choose the LogNormal distribution
$y_{n}^{k}|g,h\sim\text{LogNormal}\left(p^{g}(t_{n}),h\right),$ (12)
with an unknown scale parameter $h$. For the definition, see appendix B. Following section 2.1.1, we denote by $y_{n}^{k}$ the $k^{\text{th}}$ measurement made at time $t_{n}$ during the experiment and assume the statistics in eqs. 2 and 3 to form the reported batch mean $z^{1}_{n}$ and standard deviation $z^{2}_{n}$ of each time point.
Finally, for eq. 6, we apply a prior on the parameters $g=(Q,P,m,a)$ of our IVP that allows only for strictly positive values as defined by
$Q$ $\sim\text{Gamma}\left(\phi_{Q},\psi_{Q}/\phi_{Q}\right),$ (13)
$P$ $\sim\text{Gamma}\left(\phi_{P},\psi_{P}/\phi_{P}\right),$ (14)
$m$ $\sim\text{Gamma}\left(\phi_{m},\psi_{m}/\phi_{m}\right),$ (15)
$a$ $\sim\text{Gamma}\left(\phi_{a},\psi_{a}/\phi_{a}\right),$ (16)
@@
$h\sim\text{Gamma}\left(\phi_{h},\psi_{h}/\phi_{h}\right),$ (17)
which is conditionally conjugate to the likelihood in eq. 12. Here, we use the parameterization of the Gamma distribution, shown in appendix B, which allows the specification of shape and expectation hyperparameters via $\phi_{Q},\phi_{P},\phi_{m},\phi_{a},\phi_{h}$ and $\psi_{Q},\psi_{P},\psi_{m},\psi_{a},\psi_{h}$, respectively.
Our framework leads to the derivation of a formal posterior density
@@
Here, we marginalize the noise parameter $h$ since its value is of little biological interest. By this marginalization, we avoid the ancestral sampling steps in our MCMC scheme of section 2.1.2, but still apply mESS to update $Q,P,m,a$ and cHMC to update $y_{1:N}^{1:M}$ as previously described.